Team:Baltimore US/AugGroupNotebook


TitleBarBalti US.png
Home Team Official Team Profile Project Submitted Parts Modeling Notebook Meeting/Lab Times Safety

Group Notebook for August 2010


August 31

@ Burkett Lab Notes from Robert Buck:
Today, we have re-run the gel on the two fragments of the modified TAQ part, and the results show the
experiment to be ready to move on. We have the modified pieces and have started running a gel to purify the
products before PCRing them together. I am uploading a picture of the gel re-run (which needed re-running
due to some bad buffer the last time.) The first two lanes are 1kb and 100 bp ladder, the third is prefixed TAQ,
and the fourth is suffixed TAQ.


August 30

@ Burkett Lab WOOHOO!!! Once more into the fray...

It was a great meeting. The enthusiasm of fall semester starting filled the halls of the school, and created a warm environment to reconnect with the group after a month's downtime in a land of transport troubles.

Progress is continuing to be made on PoliColi. Some buffer problems have been addressed, and we're hoping to have successful results showing within the next week or so. We still need to ligate the ignition and then we'll need to put the whole thing into the C3 plasmid, and test it. We have had a NY iGEM team volunteer to test, but we'll need to test compared to commercial product, and folks within DIY-Bio have offered to test as well. We'll also need to show our purification protocols and write up our characterizations.

Some work on the wiki is needed. We still need pictures uploaded, and even taken of the participants. Please update days with your journal entries of what has occurred and your reflections in regards to troubleshooting so we can show a path to those without these experiences. The simplistic modeling flowcharts need to be constructed for the various parts as well.

More supplies were identified and acquired for the TraJ focus experiments being conducted at the Towson laboratory, where Liz, Lisa, Patrick and Duke will be toiling away in between classes.

Tom also discussed the ideas of working with the local high school teachers, forensic academy and DIY-Bio contingent to make up some of the kits, Miles has been working on in connection with some small experiments. Patrick and I have proposed an smaller dry run for the DIYFest coming up in October. We'll need input from Miles to see if this might be doable, in conjunction with the Harford Hackerspace contingent.

Ryan, Tom and Bernadette discussed the BioSafety Review Board and the questions iGEM has presented us with in this self-audit. Ryan also brought up his recent enrollment in ABSA and invitation to this years conference in Denver at the tail of September, suggesting it may be a perfect opportunity to show our level of commitment to proceeding cautiously and in good measure. He also reminded folks of the Public Presidential Summit on BioEthics in two weeks on the 13th & 14th in Philadelphia, to see if anyone might be interested in going.

We're awaiting permission from CCBC and TU public affairs officials still to find out whether we're clear to have the documentary folks from Germany here at the end of October that will be following us in our last few weeks up to the jamboree at MIT. Another intriguing development in media exposure.

We will continue to schedule Strategy Nights from 8-10 on Mondays at the CCBC lab (room d206) and will have workbench hours in the lab on Mondays from 12 - 5pm, as well as Tuesdays from 1 - 5pm, for those of you whom might be able to join us to work on your technique and get some of this out of the way. With this in mind, we need to keep our workflow in mind to make certain our projects will be at state that can be put away for a few days when we walk out on tuesday afternoon, since Tom's schedule is booked up with active classes throughout the week. We need to keep this in mind in regards to the equipment and supplies we're using as well, since the novices will be playing on off-hours.

This has been your BioMore Update on your Baltimore - US iGEM 2010 team....

Over and out...

August 23

@ Burkett Lab Parts to be transferred to Roberts Lab:

BBa_C00514EpSB1A2AcI repressor from E. coli phage lambda (+LVA)
BBa_R00406IpSB1A2A"TetR repressible promoter"
BBa_E04208KpSB1A2A"ECFP (RBS+ LVA- TERM) (B0034.E0020.B0015)"
BBa_J6111711LpSB1A2ARibosome Binding Site Family Member
BBa_E024012MpSB1A2AGFP generator
BBa_I050014NpSB2K3KInducible pBad/araC promoter
BBa_J2310018CBBa_J61002Aconstitutive promoter family member
BBa_I1350422IpSB1A2AScreening plasmid intermediate
BBa_J010022-22IpSB1AC3ALambda Prm Promoter

Sketch of the TraJ xo protocol:

Donor: WTF+ Kan Recipient: OriT Amp, Chlor

  1. Transform TraJ into donor
  2. Conjugate F into donor
  3. XO TraJ
  4. Select K, A, C.


  1. Conjugate WTF to recipient.(A,C)^R
  2. Recipient RecA- has OriT, inducible TraJ T^R, Select A,C,T.
  3. Scrape colonies, pool.
  4. Make competent.
  5. Transform in TraJ-Kd4 PCR product, select TACK or select K, then TAC.

Exp: see transfer of A,C iff arabinose induced.

August 16, 2010

@ Burkett's Lab Bernadette's notes:
Robert ran gels from our rxns on Aug 12th - success- Robert analyzed

We tweaked rxn conditions for optimal DNA amplification.

Bernadette's rxn primers included TAQ Rm and Bb Suufix + TAQ Reverse
PCR temps ranging from 57 degrees C to 98 for 30 cycles

Robert's rxn primers included Fwd Poll Complement and TAQ Fm
Robert ran separate PCR

August 12, 2010

@ Burkett's Lab Bernadette prepped Control for PCR (Color Coded Red) with 34 ul H20, 2ul Control Template DNA,
10 ul 5x Pfu HF Buffer, 1 ul 10mM dNTPs, 2.5 ul Primers, .5 ul Pfu DNA Polymerase

Cycling 72 - 98 degree C, hold @ 10 degrees C

Also ran Agarose gel for PCR 2..
prepped 2 gels - one refrigerated for future use
Observed apparent DNA fragment, but will re-run samples on Monday with ladders to confirm results

August 11th, 2010

@ Burkett's Lab Bernadette's notes:

PCR from Aug. 10th unsuccessful due to the low PCR temp of 72 degrees C. New primers received require low
temp of 69 degrees.
Prepared PCR 1: Rxn 1 (colored coded Black) included 2.5 ul Taq DNA, 10 ul Pfu buffer
25.5 ul H20, 1ul dNTP, 5 ul (diluted) Fwd Poll Complement, 5 ul (diluted) TAQ RM, 1 ul Pfu.
Rxn 2 (Color coded Blue) was identical to Rxn 1 except it included 1 ul DMSO & 24.5 ul H20
3rd group (Color coded Red) was a no DNA group
4th group (Color coded Green) was a no enzyme group

Ran a Touch Down PCR ranging from 61 - 98 degrees C

Prepped PCR 2:
Rxn 1 (Color coded Light Green): 2.5 ul Taq DNA, 25.5 ul H20, 1 ul dNTPs, 5 ul TAQ FM (diluted)
5 ul Bb suffix+TAQ RM (diluted), 10 ul PFU buffer & 1 ul PFU

No DNA rxn (Color Coded Orange) prepped w/ 28 ul H20 & obviously no TAQ DNA

No PFU rxn (Color Coded Brown) prepped w/ 29 ul H20 & no PFU

Touchdown PCR --cycles ranging from 61-98 degree C

Agarose gel ran for PCR 1 completed 8-11

August 10th, 2010

@ Burkett Lab Robert's notes:
New dNTPs and Primers for PCR received and aliquots made of each.
PCR optimization can be started back up once again, now that uncontaminated
materials are available again. Testing the necessity of a 2:00 minute denature time
at the start of the PCR program, as well as testing three parts
(J23056, J23031 & J23008) for DNA after boiling mini-prep from a few days ago.

For the PCR, I am using concentrations of .5 uM for primers and 200 uM for dNTPs,
a 1:1000 dilution of the three parts for boiling prep DNA, and for the other
tube's DNA, a 1:10000 dilution of J04450.

Duke ran a gel for the PCR products from last night's meeting, where he and
Bernadette had started work on the TAQ project's first step. The results were
less than he had hoped, based on his reaction to the gel.

Patrick's notes:

August 9th, 2010

@ Burkett lab Bernadette's notes:

Duke ran PCR #1

Rxn 1 consisted of 1 ul PFU, 2.5 ul Taq DNA, .5 ul FM, .5 ul RM, 10 ul PFU buffer
.5 ul dNTP & 35 ul H20.
Rxn 2 was identical to Rxn 1 except it contained 1 ul DMSO & 35 ul H20
A No DNA sample was produced and A No PFU sample was also produced

Bernadette ran PCR #2

Rxn 1 consisted of 1 ul PFU, .5 ul FM, .5 ul RM, 10 ul PFU buffer, 2.5 ul DNA, 35 ul H20
Rxn 2 was identical to 1 except that it contained 1 ul DMSO & 34 ul H20
A No DNA sample was produced (no DMSO) & a No PFU sample produced (no DMSO)

All 8 rxns ran through 30 cycles ranging from 72-98 degrees C

August 3rd, 2010

@ Burkett lab To determine validity of Castenholz media, Bernadette performed an inoculation
of Thermos Aquaticus/Taq. Due to the non antibiotic resistant Taq,
inoculation completed under sterile conditions/hood with all equipment, material
and limbs under hood cleaned with 70% isopropanol w/autoclaved pipettes.

Taq kept at -85 degrees C, heat blocked at 37 degrees C for a minute

15 ml media was combined with Taq

Incubating overnight @ 70 degrees C to observe for Taq growth.

Patrick completed DNA extractions using 23007 and 23030
Bernadette ran a gel to determine DNA presence.
DNA successfully extracted confirmed by Electrophoresis.